A chemist needs to dilute a 15% acid solution so that it contains 10% acid. How much pure water should be added to 200 milliliters of a 15% acid solution in order to reach her goal?
Answer by mananth(16946)

--- percent ---------------- Amount
Solution 15 ---------------- 200 ml
Water 0 ---------------- x ml
Mixture 10 ---------------- 200 + x ml

15 * 200 + 0 x = 10 ( 200 + x )
3000 + 0 x = 2000 + 10 x
0 x -10 x = -3000 + 2000
-10 x = -1000
/ -10
x = 100 ml Water

m.ananth@hotmail.ca